# Self-avoiding walks on cubic graphs and local transformations

### v1

@@ Abstract.
We study self-avoiding walks (SAWs) on infinite quasi-transitive cubic graphs under local transformations that replace each degree-$3$ vertex by a finite, symmetric three-port gadget. To each gadget we associate a two-port SAW generating function $g(x)$, defined by counting SAWs that enter and exit the gadget through prescribed ports. Our first main result shows that, if $G$ is cubic and $G_{1}=\phi(G)$ is obtained by applying the local transformation at every vertex, then the connective constants $\mu(G)$ and $\mu(G_{1})$ satisfy the functional relation
$\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
We next consider critical exponents defined via susceptibility-type series that do not rely on an ambient Euclidean dimension, and prove that the exponents $\gamma$ and $\eta$ are invariant under local transformations; moreover $\nu$ is invariant under a standard regularity hypothesis on SAW counts (a common slowly varying function).
Our second set of results concerns bipartite graphs, where the local transformation is applied to one colour class (or to both classes, possibly with different gadgets). In this setting we obtain an analogous relation
$\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
with $h(x)=xg(x)$ when only one class is transformed and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ when both are transformed. We further present explicit families of examples, including replacing each degree-3 vertex by a complete-graph gadget $K_{N}$.
@@ 1. Introduction
A self-avoiding walk (SAW) on a graph $G$ is a path that visits no vertex more than once. SAWs were introduced in the chemical theory of polymerization (see Flory [5]) and have since been central objects in probability, combinatorics, and statistical physics; see, for example, Madras and Slade [13]. For an infinite quasi-transitive graph $G$, the number $\sigma_{n}$ of $n$-step SAWs grows exponentially, and the associated exponential growth rate
$\mu=\mu(G)=\lim_{n\to\infty}\sigma_{n}^{1/n}$
@@
is the connective constant of $G$. Exact values of connective constants are known only in a small number of nontrivial cases.
This paper studies how SAWs and their asymptotic parameters behave under local transformations of cubic graphs. Informally, a local transformation replaces every degree-$3$ vertex by a fixed finite gadget with three distinguished outer vertices (ports) attached to the three incident edges, subject to a symmetry condition ensuring that the three ports are indistinguishable from the point of view of the gadget (Definition 3.1). A canonical example is the Fisher transformation, which replaces each vertex by a triangle; Fisher-type transformations play an important role in exact mappings between two-dimensional lattice models (e.g. Ising/dimer/vertex models) and have also been used in the SAW literature.
@@
Connective constants under local transformations. Given a local transformation $\phi$, we associate to the gadget a two-port generating function $g(x)$ (Definition 3.2), defined by counting SAWs that enter the gadget through one port-edge and exit through another, while otherwise staying inside the gadget. Our first main theorem (Theorem 3.3) shows that this local data determines the connective constant of the transformed graph: if $G$ is infinite, connected, quasi-transitive and cubic, and $G_{1}=\phi(G)$ is obtained by applying $\phi$ at every vertex, then
(1.1) $\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
Equivalently, $\mu(G_{1})^{-1}$ is the unique positive solution of $g(x)=\mu(G)^{-1}$. This functional relation may be viewed as a substitution identity for SAW partition functions, and it extends the familiar relation arising from the Fisher transformation.
@@
Critical exponents. Beyond exponential growth, SAWs are expected to exhibit universal power-law corrections, captured by critical exponents. Rather than assuming an ambient Euclidean dimension, we work with graph-theoretic definitions of the exponents $\gamma,\eta,\nu$ via the finiteness of natural SAW generating series (Section 3). Our second main theorem (Theorem 3.4) proves that local transformations preserve these exponents under mild hypotheses: $\gamma$ and $\eta$ are invariant, and $\nu$ is invariant assuming the SAW counts satisfy a standard slowly varying correction (3.7) with a common slowly varying function.
Transforming one colour class of a bipartite graph. Our third main result concerns bipartite graphs. Let $G$ be an infinite, connected, quasi-transitive bipartite graph with colour classes (black/white), and suppose at least one class is cubic. Let $G_{\mathrm{e}}$ be obtained by applying a local transformation to all vertices in one class (or, in the fully cubic case, applying possibly different local transformations to the two classes). Theorem 3.5 shows that the connective constants satisfy
(1.2) $\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
with $h(x)=xg(x)$ if only one class is transformed, and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ if both classes are transformed. This yields new exact connective constants when combined with known exact values.
The special case in which each cubic vertex is replaced by a triangle is the Fisher transformation. The effect of the Fisher transformation on self-avoiding walks was analyzed by Grimmett and Li [6], who derived an explicit functional relation between the connective constants of $G$ and its Fisher transform and proved invariance of the principal critical exponents under iteration.
The present paper extends [6] from Fisher triangles to general local transformations in which each degree-$3$ vertex is replaced by an arbitrary finite three-port gadget satisfying a natural symmetry condition (Definition 3.1). Under this general framework we obtain corresponding functional relations for connective constants (Theorems 3.3 and 3.5) and prove invariance of the critical exponents under verifiable hypotheses (Theorem 3.4), including the case where the transformation is applied to only one colour class of a bipartite graph (Theorem 3.5). .
Further examples. We include explicit families of gadgets beyond Fisher triangles. In particular, replacing each cubic vertex by a complete graph $K_{N}$ with three designated outer vertices yields an explicit polynomial $g_{N}(x)$ (Example 6.1), and hence an explicit implicit relation between the connective constants of $G$ and $\phi^{(N)}(G)$ for any $N\geq 3$. When $N=3$, this corresponds to Fisher transformation. See Figure 1.1 for local replacements of a degree-3 vertex by a complete graph $K_{3}$ and $K_{4}$.
[…]
Figure $1.$ Local replacement at a degree-$3$ vertex: (a) the original star, (b) Fisher transformation (triangle gadget), and (c) replacement by a $K_{4}$ gadget. See also Definition 3.1 and Figure 3.1 in the draft.
A useful class of examples comes from vertex-transitive TLF-planar cubic graphs. Following Renault ([14]), planarity plus vertex-transitivity imposes strong geometric symmetry on embeddings: the action of graph automorphisms can (locally) be represented by isometries of the ambient geometry (spherical, Euclidean, or hyperbolic).
Let $G$ be a transitive, TLF-planar cubic graph, and let $\langle a,b,c\rangle$ be its type-vector, i.e., the sizes of the three faces incident to a vertex, with the convention $1/\infty=0$. Define
$f(G):=\frac{1}{a}+\frac{1}{b}+\frac{1}{c}.$
By [14] (see also [8]), $G$ admits a planar TLF embedding by regular polygons in the sphere / Euclidean plane / hyperbolic plane according as $f(G)>\frac{1}{2}$, $f(G)=\frac{1}{2}$, or $f(G)<\frac{1}{2}$,
In the Euclidean case $f(G)=\frac{1}{2}$, the only possible type-vectors are $\langle 6,6,6\rangle$, $\langle 3,12,12\rangle$, $\langle 4,8,8\rangle$, $\langle 4,6,12\rangle$, and $\langle 4,4,\infty\rangle$; see Figure 1.2.
$\langle 6,6,6\rangle$ $\langle 4,8,8\rangle$ $\langle 3,12,12\rangle$
$\langle 4,6,12\rangle$ $\langle 4,4,\infty\rangle$
Figure 1.2. Local regular-polygon patches for the five Euclidean type-vectors of transitive TLF-planar cubic graphs.
Organization. Section 2 collects basic notation. Local transformations and the associated generating function $g(x)$ are introduced in Section 3, where we also state Theorems 3.3–3.5. The proofs of Theorems 3.3–3.4 and Theorem 3.5 appear in Sections 4 and 5, respectively. We conclude with explicit examples in Section 6.
@@ 3. Local transformation
where $\sigma_{n}(v,w)$ is the number of $n$-step SAWs with endpoints $v$, $w$. It is known under certain circumstances that the generating functions $Z_{v,w}$ have radius of convergence $\mu^{-1}$ (see [13, Cor. 3.2.6]), and it is believed that there exists an exponent $\eta$ and constants $A^{\prime}_{v}>0$ such that
@@
6. Example
Example 6.1 (Complete-graph gadget $K_{N}$).
@@ 6. Example
By Theorem 3.3, if $G$ is cubic with connective constant $\mu$ and $G_{1}=\phi^{(N)}(G)$ has connective constant $\mu_{1}$, then $\mu^{-1}=g_{N}(\mu_{1}^{-1})$. In particular, $\mu_{1}^{-1}$ is the unique solution $x\in(0,1)$ of $g_{N}(x)=\mu^{-1}$.
@@
[…]
Figure 6.2. A generalized Fisher transformation: (a) a cycle graph $H$ on $N=7$ vertices with two distinguished vertices $v_{1},v_{2}$ spaced $k=2$ edges apart and (b) the gadget $\widetilde{H}$ constructed according to Example 6.2.

### v2

@@ Abstract.
Despite its elementary definition, the self-avoiding walk (SAW) poses notoriously hard enumerative problems: exact connective constants are known for only a handful of infinite graphs, notably the honeycomb lattice [3]. We establish a general substitution principle for SAWs on infinite connected quasi-transitive cubic graphs under port-transitive vertex replacements, where each degree-$3$ vertex is replaced by a fixed finite three-port gadget. Writing $g(x)$ for the associated two-port SAW series, we prove that for $G_{1}=\phi(G)$,
$\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr),$
equivalently $\mu(G_{1})^{-1}$ is the unique solution $x\in(0,1)$ of $g(x)=\mu(G)^{-1}$, thereby extending the Fisher-triangle relation of Grimmett–Li to arbitrary symmetric three-port gadgets. We also obtain the corresponding identity for bipartite graphs when one or both colour classes are transformed, and show that the critical exponents $\gamma$ and $\eta$ (and $\nu$ under a standard regularity hypothesis) are invariant. For explicit gadget families, including complete-graph gadgets $K_{N}$ and Fisher-type constructions, these identities turn base graphs with known $\mu$ into infinite families of new quasi-transitive graphs whose connective constants are determined exactly as the unique roots of explicit algebraic equations.
@@ 1. Introduction
A self-avoiding walk (SAW) on a graph is a finite path that visits no vertex more than once. Introduced as a model for polymer chains, SAWs have become central objects in probability, combinatorics, and statistical physics; see, for example, [5, 12]. Let $G$ be an infinite connected quasi-transitive graph, and let $\sigma_{n}(v)$ denote the number of $n$-step SAWs starting at $v$. A classical theorem of Hammersley implies that the limit
$\mu(G)=\lim_{n\to\infty}\sigma_{n}(v)^{1/n}$
@@
exists and is independent of $v$; the constant $\mu(G)$ is the connective constant of $G$. Exact values (or explicit characterizations) of $\mu(G)$ are known only in a small number of nontrivial cases.
Local vertex replacements and a substitution principle. We study SAWs on cubic graphs under a class of vertex-replacement operations. A local transformation $\phi$ replaces each degree-$3$ vertex by a fixed finite three-port gadget whose three ports attach to the incident edges, subject to a port-transitivity condition on the gadget automorphism group (Definition 3.1). The classical example is the Fisher transformation, where each vertex is replaced by a triangle.
Grimmett and Li [6] analysed SAWs under the Fisher transformation, obtaining an explicit functional relation between connective constants and proving invariance of key critical exponents under iterated Fisher transforms. The aim of the present paper is to establish a general exact substitution framework that (i) extends the Fisher-triangle analysis to arbitrary port-transitive three-port gadgets, and (ii) treats the bipartite setting in which the transformation is applied to one colour class, or to both classes with possibly different gadgets. The resulting identities are global (they describe radii of convergence and critical exponents) but are driven by finite, explicitly computable local data.
A two-port gadget series and connective constants. Fix $\phi$, and let $H$ denote its gadget. We associate to $H$ a two-port generating function
(1.1) $g(x)=\sum_{\pi}x^{|\pi|},$
@@
where $\pi$ ranges over SAWs that enter $H$ through one prescribed port-edge and exit through another while otherwise remaining inside $H$ (Definition 3.2). Our first main theorem (Theorem 3.3) shows that this purely local series governs the connective constant after transforming every vertex: if $G$ is infinite, connected, quasi-transitive, and cubic, and $G_{1}=\phi(G)$, then
(1.2) $\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr),$
equivalently $\mu(G_{1})^{-1}$ is the unique $x\in(0,1)$ solving $g(x)=\mu(G)^{-1}$.
@@
Critical exponents on quasi-transitive graphs. To avoid assuming an ambient Euclidean dimension, we adopt graph-theoretic definitions of critical exponents $\gamma,\eta,\nu$ via finiteness of natural susceptibility-type series (Section 3). Our second main theorem (Theorem 3.4) proves that local transformations preserve these exponents under mild hypotheses: $\gamma$ and $\eta$ are invariant in general, and $\nu$ is invariant provided the SAW counts satisfy a standard regularity hypothesis with a common slowly varying correction.
Bipartite transformations. If $G$ is bipartite and the transformation is applied to one colour class (or, in the fully cubic case, to both classes with possibly different gadgets), we obtain an analogous relation (Theorem 3.5):
$\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
with $h(x)=xg(x)$ in the one-class case and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ in the two-class case. These formulas turn known values of $\mu(G)$ into explicit equations for $\mu(G_{\mathrm{e}})$.
Examples and computations. Section 6 develops explicit gadget families and worked base-graph examples. We emphasize complete-graph gadgets $K_{N}$ (yielding closed-form polynomials $g_{N}$), a mixed bipartite transformation on the hexagonal lattice (white $K_{4}$, black Fisher triangles), and an iterated transformation example that exhibits convergence of the sequence of connective constants to a gadget-determined fixed point.
Organization. Section 2 introduces notation. Section 3 defines local transformations and the associated series $g(x)$ and states Theorems 3.33.5. Proofs appear in Sections 4 and 5. Section 6 contains explicit examples, computations, and remarks on iteration.
@@ 3. Local transformation
where $\sigma_{n}(v,w)$ is the number of $n$-step SAWs with endpoints $v$, $w$. It is known under certain circumstances that the generating functions $Z_{v,w}$ have radius of convergence $\mu^{-1}$ (see [12, Cor. 3.2.6]), and it is believed that there exists an exponent $\eta$ and constants $A^{\prime}_{v}>0$ such that
@@
6. Examples
This section collects explicit gadget families and worked base-graph examples illustrating Theorems 3.3 and 3.5. Many natural quasi-transitive cubic graphs are planar (e.g. the hexagonal lattice), while others are not (e.g. the $3$-regular tree); our substitution identities apply uniformly in all cases.
Example 6.1 (Complete-graph gadgets $K_{N}$).
@@ 6. Examples
By Theorem 3.3, if $G$ is cubic with connective constant $\mu$ and $G_{1}=\phi^{(N)}(G)$ has connective constant $\mu_{1}$, then
$\mu^{-1}=g_{N}(\mu_{1}^{-1}),$
and $\mu_{1}^{-1}$ is the unique solution $x\in(0,1)$ of $g_{N}(x)=\mu^{-1}$.
@@
Remark 6.2 (A worked base graph: the $3$-regular tree).
Let $T_{3}$ denote the $3$-regular tree. Then $\mu(T_{3})=2$ (indeed $\sigma_{n}=3\cdot 2^{n-1}$). For $G_{1}=\phi^{(N)}(T_{3})$, Theorem 3.3 yields $g_{N}(\mu_{1}^{-1})=\tfrac{1}{2}$. For instance,
$\mu_{1}\approx 2.16730\ (N=4),\qquad\mu_{1}\approx 2.60068\ (N=5),\qquad\mu_{1}\approx 3.55052\ (N=7),$
obtained by solving $g_{N}(x)=\tfrac{1}{2}$ for $x\in(0,1)$ and setting $\mu_{1}=1/x$. This illustrates how complete-graph gadgets produce explicit, finite polynomial equations for new connective constants.
@@
Example 6.3 (A planar six-vertex “triangle-in-triangle” gadget).
Let $\phi^{\triangle}$ be the local transformation whose gadget $T$ is the following $6$-vertex graph (see Figure 6.2). The gadget has three ports (outer vertices) $w_{1},w_{2},w_{3}$. On each side $w_{i}w_{j}$ of the outer triangle we insert one vertex $u_{ij}$, so the boundary edges are
$w_{1}-u_{12}-w_{2},\qquad w_{2}-u_{23}-w_{3},\qquad w_{3}-u_{31}-w_{1},$
and we add the three edges forming the inner triangle
$u_{12}u_{23},\qquad u_{23}u_{31},\qquad u_{31}u_{12}.$
The dihedral symmetries act transitively on $\{w_{1},w_{2},w_{3}\}$, so $\phi^{\triangle}$ satisfies Definition 3.1.
Let $g_{\triangle}(x):=g_{\phi^{\triangle}}(x)$ be defined by (3.1). A gadget-SAW from the port-edge at $w_{1}$ to that at $w_{2}$ is exactly a simple path from $w_{1}$ to $w_{2}$ inside $T$ (mid-edge convention). Enumerating simple paths by their number of visited vertices gives
$g_{\triangle}(x)=x^{3}+3x^{4}+4x^{5}+2x^{6}.$
Therefore, for any infinite connected quasi-transitive cubic graph $G$ with connective constant $\mu$ and $G_{1}=\phi^{\triangle}(G)$ with connective constant $\mu_{1}$, Theorem 3.3 yields the exact relation
$\mu^{-1}=g_{\triangle}\!\bigl(\mu_{1}^{-1}\bigr).$
Hexagonal lattice. If $G=\mathbb{H}$ is the hexagonal lattice, then $\mu(\mathbb{H})=\sqrt{2+\sqrt{2}}$ [3], so $\mu(\phi^{\triangle}(\mathbb{H}))^{-1}\in(0,1)$ is the unique solution of
$x^{3}+3x^{4}+4x^{5}+2x^{6}=\bigl(2+\sqrt{2}\bigr)^{-1/2}.$
(Numerically, this gives $\mu(\phi^{\triangle}(\mathbb{H}))\approx 1.9318771$.)
AABBCCtriangle-in-triangle gadget
Figure 6.2. A planar 66-vertex gadget with outer ports $w_{1},w_{2},w_{3}$ and an inner triangle supported on the three sides. For this gadget, $g_{\triangle}(x)=x^{3}+3x^{4}+4x^{5}+2x^{6}$.
Example 6.4 (A mixed bipartite transformation on the hexagonal lattice).
Let $H$ be the hexagonal lattice with its natural bipartition $V_{\mathrm{w}}\cup V_{\mathrm{b}}$ (white/black). Apply $\phi^{(4)}$ at each white vertex (a $K_{4}$ gadget), and apply the Fisher triangle transformation at each black vertex (i.e. a $K_{3}$ gadget). Denote the resulting graph by $H_{\mathrm{e}}$.
For the Fisher triangle one has
$g_{K_{3}}(x)=x^{2}+x^{3},$
and for the $K_{4}$ gadget one has $g_{K_{4}}(x)=x^{2}+2x^{3}+2x^{4}$. By Theorem 3.5(2),
$\mu(H)^{-2}=g_{K_{4}}\bigl(\mu(H_{\mathrm{e}})^{-1}\bigr)\,g_{K_{3}}\bigl(\mu(H_{\mathrm{e}})^{-1}\bigr).$
Since $\mu(H)^{-2}=(2+\sqrt{2})^{-1}=\tfrac{1}{2}(2-\sqrt{2})$, writing $x:=\mu(H_{\mathrm{e}})^{-1}$ yields the explicit polynomial equation
(6.2) $x^{4}+3x^{5}+4x^{6}+2x^{7}=\tfrac{1}{2}(2-\sqrt{2}),\qquad x\in(0,1).$
Thus $\mu(H_{\mathrm{e}})=1/x\approx 1.91529$.
The two gadgets used here are depicted in Figure 6.3.
[…]
Figure 6.3. The two gadgets used in Example 6.4: a Fisher triangle ($K_{3}$) and a $K_{4}$ gadget with ports $A,B,C$.
@@
Example 6.7 (A “hub–triangle” gadget and an explicit iteration picture).
Define a local transformation $\psi$ by replacing each degree-$3$ vertex by the gadget obtained as follows: start from a triangle on outer vertices $w_{1},w_{2},w_{3}$ (the ports), subdivide each of the three triangle edges by inserting a new vertex at its midpoint, and finally add a central vertex connected to the three midpoint vertices. (See Figure 6.4.) This gadget is port-transitive by the dihedral symmetry of the construction.
A direct enumeration of gadget SAWs (in the mid-edge convention of Section 2) yields
(6.4) $g_{\psi}(x)=x^{3}+4x^{5}+2x^{7}.$
In particular, $g_{\psi}$ is strictly increasing and strictly convex on $(0,1)$ and satisfies $g_{\psi}(1)=7$.
If $G_{n}:=\psi^{n}(G)$ remains cubic for all $n\geq 0$ and $x_{n}:=\mu(G_{n})^{-1}$, then Theorem 3.3 gives $x_{n-1}=g_{\psi}(x_{n})$ for $n\geq 1$. Since $g_{\psi}$ is strictly convex with $g_{\psi}(0)=0$ and $g_{\psi}^{\prime}(0)=0$, the equation $g_{\psi}(x)=x$ has a unique solution $x_{\ast}\in(0,1)$, and the sequence $(x_{n})$ is monotone and converges to $x_{\ast}$ (cf. Remark 6.8 below). Numerically, $x_{\ast}\approx 0.6050003$, hence the limiting connective constant is $1/x_{\ast}\approx 1.65289$.
    AABBCC(a) triangle–hub gadget     $yxx_{\ast}$(b) $y=g_{\psi}(x)$ and $y=x$
Figure 6.4. (a) The triangle–hub gadget used to define the local transformation $\psi$. (b) The fixed point $x_{\ast}$ solving $g_{\psi}(x)=x$ for $g_{\psi}(x)=x^{3}+4x^{5}+2x^{7}$. So $x_{\ast}=\sqrt{\frac{\sqrt{3}-1}{2}}\approx 0.6050003$, and $\mu(\psi^{n}(G))\longrightarrow\mu_{\ast}:=x_{\ast}^{-1}\approx 1.6528917$.
@@
Acknowledgements
The authors thank the organizers of the conference “Perfectly matched perspectives on statistical mechanics, combinatorics and geometry” (CIRM, 16–20 June 2025) for a stimulating meeting. The conference was supported by the National Science Foundation under Grant DMS-2512154. B.G. was partially supported by a U.S. Department of Education GAANN Fellowship.